A system undergoes a two-step process. In the first step, the internal energy of the system increases by 226 J when 166 J of work is done on the system. In the second step, the internal energy of the system increases by 115 J when 165 J of work is done on the system. For the overall process, find the heat added to the system (including sign).
Answer #1

As E=q+w ==> q=E-w

Heat added to the system=(226-166)J + (115-165) J =10 J